OpenStudy (anonymous):

Guts1851: Sorry in this Find the slope and y-intercept for the given equation. x - y = 2 m = b = Guts1851: Sorry in this Find the slope and y-intercept for the given equation. x - y = 2 m = b = @Mathematics

OpenStudy (anonymous):

m = 1 b = -2

OpenStudy (anonymous):

Find the slope of the following line. The line 5 x - 3 y = 4 m =

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

slope = 5/3 intercept = -4/3